# Arbitrage-free catastrophe reinsurance valuation for compound dynamic contagion claims

### v1

@@ 5 Conclusion
Unless (re)insurers have an appropriate framework and prudent methodology which adequately assess catastrophe risks, ordinary individuals and business owners will suffer as the industry will not be able to promptly and fully meet the many claims of catastrophe victims. To address this challenge, in this paper, we presented how to obtain arbitrage-free catastrophe stop-loss reinsurance premiums when dealing with the ongoing challenge and complexity of emerging risk dynamics. To predict claim/loss arrivals from conventional and emerging catastrophes, we used the compound dynamic contagion process for the catastrophic component of the liability and the Esscher transform to determine arbitrage-free premiums. Sensitivity analyses were also performed by changing the Esscher parameters and the retention level. Given the broad applicability of the CDCP, it is expected that what we have obtained in this paper will provide practitioners with feasible approaches to quantify their catastrophic liabilities in light of the growing challenges posed by new risks arising from climate change, cyberattacks, and pandemics. We also envisage that the dynamics used in this paper extend naturally for pricing a variety of catastrophe insurance derivatives. For this, we leave them as objects of further research.
Data
The Python code used to generate the numerical results is available at:
https://github.com/Pat-Laub/ReinsuranceEsscherDynamicContagion
Appendix A Simulation algorithms
Algorithm A.2 shows our method for simulating the time-inhomogenous compound dynamic contagion process, which extends the work of Ogata (1981). It utilises Algorithm A.1 which is the well-known thinning algorithm for Poisson processes (see Lewis and Shedler, 1979).
Algorithm A.1 Simulate a time-inhomogeneous Poisson process
1:procedure SampleInhomogeneousPoissonProcess($T,\;\rho(\cdot)$)
@@ Appendix A Simulation algorithms
4:     Draw $N$ candidate times $\{t^{*}_{1},\ldots,t^{*}_{N}\}$ uniformly in $[0,T]$ and sort them
5:     Initialize $\mathcal{T}\leftarrow\emptyset$
@@
In the next algorithm we split the intensity into
$\lambda_{t}=\underbrace{a(t)+\left(\lambda_{0}-a(t)\right)\mathrm{e}^{-\delta t}+\sum_{i\geq 1}X_{i}\mathrm{e}^{-\delta\left(t-T_{1,i}\right)}\mathbb{I}_{\{T_{1,i}\leq t\}}}_{=:\lambda^{(c)}_{t},\text{ the ``Cox'' part}}+\underbrace{\sum_{j\geq 1}Y_{j}\mathrm{e}^{-\delta\left(t-T_{2,j}\right)}\mathbb{I}_{\{T_{2,j}\leq t\}}}_{=:\lambda^{(h)}_{t},\text{ the ``Hawkes'' part}}\,.$
Algorithm A.2 Simulate a time-inhomogeneous compound dynamic contagion process
1:procedure SimulateCDCP($\mathcal{T}=[0,T],\;\lambda_{0},\;a(t),\;\rho(t),\;\delta,\;G(\cdot;t),\;H(\cdot;t),\;J(\cdot),\;\Delta t_{\text{max}}$)
2:     Let $\{t_{1,1},t_{1,2},\ldots,t_{1,n}\}\leftarrow\textsc{SampleInhomogeneousPoissonProcess}(T,\;\rho(\cdot))$
3:     for all $i=1,2,\ldots,n$ do
4:         Sample $X_{i}\sim H(\,\cdot\,;t_{1,i})$
5:     Compute $\lambda^{(c)}_{t}\leftarrow a(t)+(\lambda_{0}-a(t))\mathrm{e}^{-\delta t}+\sum_{i:t_{1,i}\leq t}X_{i}\,\mathrm{e}^{-\delta(t-t_{1,i})}$ for all $t\in\mathcal{T}$
6:     Initialize $\lambda^{(h)}_{t}\leftarrow 0$ for all $t\in\mathcal{T}$, $t\leftarrow 0$ and $j\leftarrow 0$
7:     while $t\leq T$ do
8:         Compute $\lambda_{\max}\leftarrow\max_{t^{\prime}\in[t,(t+\Delta t_{\text{max}})\wedge T]}\ \lambda^{(c)}_{t^{\prime}}+\lambda^{(h)}_{t^{\prime}}$
9:         Sample $\Delta t\sim\mathrm{Exponential}(\lambda_{\max})$
10:         Update $t\leftarrow t+(\Delta t\wedge\Delta t_{\text{max}})$
11:         if $t>T$ then break
12:         if $\Delta t>\Delta t_{\text{max}}$ then continue
13:         Sample $U\sim\mathrm{Uniform}(0,1)$
14:         if $U<(\lambda^{(c)}_{t}+\lambda^{(h)}_{t})/\lambda_{\max}$ then
15:              $j\leftarrow j+1$ and $t_{2,j}\leftarrow t$
16:              Sample $Y_{j}\sim G(\,\cdot\,;t_{2,j})$triangleright Next Hawkes intensity jump
17:              Update $\lambda^{(h)}_{t^{\prime}}\leftarrow\lambda^{(h)}(t^{\prime})+Y_{j}\,\mathrm{e}^{-\delta(t^{\prime}-t_{2,j})}$ for all $t^{\prime}\in[t_{2,j},T]$
18:              Sample $\Xi_{j}\sim J(\,\cdot\,)$triangleright Next jump in the CDCP process
19:     return $C_{t}=\sum_{j\geq 1}\Xi_{j}\,\mathbb{I}_{\{t_{2,j}\leq t\}}$ for all $t\in\mathcal{T}$

### v2

@@ 5 Conclusion
Unless (re)insurers have an appropriate framework and prudent methodology which adequately assess catastrophe risks, ordinary individuals and business owners will suffer as the industry will not be able to promptly and fully meet the many claims of catastrophe victims. To address this challenge, in this paper, we presented how to obtain arbitrage-free catastrophe stop-loss reinsurance premiums when dealing with the ongoing challenge and complexity of emerging risk dynamics. To predict claim/loss arrivals from conventional and emerging catastrophes, we used the compound dynamic contagion process for the catastrophic component of the liability and the Esscher transform to determine arbitrage-free premiums. Sensitivity analyses were also performed by changing the retention level, the Esscher parameters and the intensity parameters. Given the broad applicability of the CDCP, it is expected that what we have obtained in this paper will provide practitioners with feasible approaches to quantify their catastrophic liabilities in light of the growing challenges posed by new risks arising from climate change, cyberattacks, and pandemics. We also envisage that the dynamics used in this paper extend naturally for pricing a variety of catastrophe insurance derivatives. For this, we leave them as objects of further research.
Competing Interests
The authors declare none.
Data Availability Statement
The Python code used to generate the numerical results is available at: https://github.com/Pat-Laub/ReinsuranceEsscherDynamicContagion.
Funding Statement
This work received no specific grant from any funding agency, commercial or not-for-profit sectors.
Appendix A Equivalent martingale measure with non-zero interest rate
Definition A.1 (Equivalent martingale measure with discounting).
Let $r\geq 0$ be a constant risk-free interest rate. A probability measure $\tilde{\mathbb{P}}$ is said to be an equivalent martingale probability measure (i.e., $\tilde{\mathbb{P}}$ is equivalent to $\mathbb{P}$ on ${\cal F}_{T}$) if it satisfies the following properties:
  1. (i)
    $\tilde{\mathbb{P}}(A)=0$ if and only if $\mathbb{P}(A)=0$ for any $A\in{\cal F}_{T}$;
  2. (ii)
    The Radon–Nikodym derivative $\frac{\mathop{}\!\mathrm{d}\tilde{\mathbb{P}}}{\mathop{}\!\mathrm{d}\mathbb{P}}\in L^{2}(\Omega,{\cal F}_{T},\mathbb{P})$;
  3. (iii)
[…]
    $\tilde{\mathbb{E}}\!\left[\mathrm{e}^{-rt}R_{t}\mid{\cal F}_{s}\right]=\mathrm{e}^{-rs}R_{s},\qquad 0\leq s\leq t\leq T,\quad\tilde{\mathbb{P}}\text{-a.s.}$
Remark A.1.
Assuming $R_{0}=0$, Definition A.1 (iii) implies
$\tilde{\mathbb{E}}\!\left(\mathrm{e}^{-rt}R_{t}\mid\mathcal{F}_{0}\right)=\tilde{\mathbb{E}}\!\left(\mathrm{e}^{-rt}(P_{t}-C_{t})\mid\mathcal{F}_{0}\right)=0.$
Consequently,
$\tilde{\mathbb{E}}\!\left(\mathrm{e}^{-rt}C_{t}\mid\mathcal{F}_{0}\right)=\tilde{\mathbb{E}}\!\left(\mathrm{e}^{-rt}P_{t}\mid\mathcal{F}_{0}\right),$
which characterises the arbitrage-free insurance premium under a non-zero constant interest rate.
Appendix B Simulation algorithms
Algorithm B.2 shows our method for simulating the time-inhomogeneous compound dynamic contagion process, which extends the work of Ogata (1981). It utilises Algorithm B.1 which is the well-known thinning algorithm for Poisson processes (see Lewis and Shedler, 1979).
Algorithm B.1 Simulate a time-inhomogeneous Poisson process
[…]
@@ Appendix B Simulation algorithms
4:  Draw NN candidate times $\{t^{*}_{1},\ldots,t^{*}_{N}\}$ uniformly in $[0,T]$ and sort them
5:  Initialise $\mathcal{T}\leftarrow\emptyset$
@@
In the next algorithm we split the terms in Eq. 2.5 into
$\lambda_{t}=\underbrace{\lambda_{0}\mathrm{e}^{-\delta t}+\delta\int_{0}^{t}a(s)\mathrm{e}^{-\delta(t-s)}\,\mathop{}\!\mathrm{d}s+\sum_{i\geq 1}X_{i}\mathrm{e}^{-\delta\left(t-T_{1,i}\right)}\mathbb{I}_{\{T_{1,i}\leq t\}}}_{=:\lambda^{(c)}_{t},\text{ the ``Cox'' part}}+\underbrace{\sum_{j\geq 1}Y_{j}\mathrm{e}^{-\delta\left(t-T_{2,j}\right)}\mathbb{I}_{\{T_{2,j}\leq t\}}}_{=:\lambda^{(h)}_{t},\text{ the ``Hawkes'' part}}\,.$ (B.1)
Algorithm B.2 Simulate a time-inhomogeneous compound dynamic contagion process
[…]
2:
3:  // Step 1: Sample External (Cox) Arrival Times and Jump Sizes
4:  Let $\{t_{1,1},t_{1,2},\ldots,t_{1,n}\}\leftarrow\textsc{SampleInhomogeneousPoissonProcess}(T,\;\rho(\cdot))$
5:  for all $i=1,2,\ldots,n$ do
6:   Sample $X_{i}\sim H(\,\cdot\,;t_{1,i})$
7:
8:  // Step 2: Pre-compute the Cox Component of the Intensity
9:  Compute $\lambda^{(c)}_{t}\leftarrow\lambda_{0}\mathrm{e}^{-\delta t}+\delta\int_{0}^{t}a(s)\mathrm{e}^{-\delta(t-s)}\,\mathop{}\!\mathrm{d}s+\sum_{i\geq 1}X_{i}\mathrm{e}^{-\delta\left(t-T_{1,i}\right)}\mathbb{I}_{\{T_{1,i}\leq t\}}$ for all $t\in\mathcal{T}$
10:
11:  // Step 3: Simulate Self-Excited (Hawkes) Arrivals via Ogata’s Thinning
12:  Initialise $\lambda^{(h)}_{t}\leftarrow 0$ for all $t\in\mathcal{T}$, $t\leftarrow 0$ and $j\leftarrow 0$ $\triangleright$ Note that $\lambda_{t}=\lambda^{(c)}_{t}+\lambda^{(h)}_{t}$
13:
14:  Set $t\leftarrow 0$ and $j\leftarrow 0$
15:  while $t\leq T$ do
16:   // Generate a potential arrival time
17:   Compute $\lambda_{\max}\leftarrow\max_{t^{\prime}\in[t,(t+\Delta t_{\text{max}})\wedge T]}\ \lambda^{(c)}_{t^{\prime}}+\lambda^{(h)}_{t^{\prime}}$
18:   Sample $\Delta t\sim\mathrm{Exponential}(\lambda_{\max})$
19:   Update $t\leftarrow t+(\Delta t\wedge\Delta t_{\text{max}})$
20:
21:   // Stop simulating if we reach the end of the simulation horizon
22:   if $t>T$ then
23:     break
24:
25:   // Discard proposals that are too far in the future
26:   if $\Delta t>\Delta t_{\text{max}}$ then
27:     continue
28:
29:   // Keep the proposal with probability $\lambda_{t}/\lambda_{\text{max}}$
30:   Sample $U\sim\mathrm{Uniform}(0,1)$
31:   if $U<(\lambda^{(c)}_{t}+\lambda^{(h)}_{t})/\lambda_{\max}$ then
32:     // Update various counters & intensities and sample a jump size
33:     $j\leftarrow j+1$ and $t_{2,j}\leftarrow t$
34:     Sample $Y_{j}\sim G(\,\cdot\,;t_{2,j})$ $\triangleright$ Next Hawkes intensity jump
35:     Update $\lambda^{(h)}_{t^{\prime}}\leftarrow\lambda^{(h)}_{t^{\prime}}+Y_{j}\,\mathrm{e}^{-\delta(t^{\prime}-t_{2,j})}$ for all $t^{\prime}\in[t_{2,j},T]$
36:     Sample $\Xi_{j}\sim J(\,\cdot\,)$ $\triangleright$ Next jump in the CDCP
37:
38:  return $C_{t}=\sum_{j\geq 1}\Xi_{j}\,\mathbb{I}_{\{t_{2,j}\leq t\}}$ for all $t\in\mathcal{T}$